OpenStudy (gigi1103):

I WILL FAN AND MEDAL Simplify. 3.5 × 22 + 12 ÷ 4 - (1.4 + 4.9) 17.3 20.5 10.7 20.7 FOR 10 QUESTIONS

OpenStudy (anonymous):

(1.4 + 4.9 = 6.3) 3.5 x 22 = 77 + 12 = 89 / divided by 4 = 22.25 - 6.3 = 15.95 so I say it would be 17.3 because its the closest to what I got